OpenStudy (anonymous):

Solve 15x^4-28x^2+5=0

OpenStudy (goformit100):

fatorize it to get the value of x.

OpenStudy (saifoo.khan):

To make things easier, x^2 = y 15y^2 - 28y + 5 = 0

OpenStudy (anonymous):

I'm confused im sorry...my brained is fried

OpenStudy (goformit100):

@saifoo.khan

OpenStudy (saifoo.khan):

If you wanted to try you may proceed, i didn't wanted to interrupt. :) @goformit100

OpenStudy (anonymous):

saifooo! did you switch with lbg whatever?

OpenStudy (anonymous):

factor as \((3 x^2-5) (5 x^2-1) = 0\) and work from there. someone is looking a lot like saifoo these days

OpenStudy (saifoo.khan):

Dang! my PC froze.

OpenStudy (anonymous):

Ok ;-)

OpenStudy (anonymous):

freezing my pc off here too

OpenStudy (saifoo.khan):

@satellite73 , People are following my footsteps to be on 99. :D ANd im following your steps to get on 100. :D

OpenStudy (saifoo.khan):

shh.

OpenStudy (saifoo.khan):

;)

OpenStudy (anonymous):

@alisa845 you got this? set each factor = 0, solve for \(x^2\) and then solve for \(x\)

OpenStudy (anonymous):

15x^2-25x^2-5

OpenStudy (anonymous):

Is that correct?